Suppose (X,Y)∼Unif(A), whereA⊆R2 is the triangular region connecting vertices (0,2),(2,0), and (2,2). FindCor(X,Y).
Answer #1

ly (x, y) = a (x,y) A II. area (A) OLX22, 044 22 x+y=2 y Tylwy dy a area (A)=LX 2X2=2 (290) 2 2-X SI dy = 04442 due to symmet